Kb F3 (s pts) Calculate the pll of the following three-component mixture: 30.0 mL of 0.200 M ethylamine (C2HsNH2 20.0 mL of 0.400 M ethyl amine hydrochloride (C2HsNHC) 50.0 ml of 0.120 M NaoH Chis is just one mixture with all three of the components above combined, so calculate just one pH for the whole thing, NOT three separate pH values!)
Answer #1

On mixing these three components we would have acid base reaction between salt C2H5NH3Cl and NaOH. This reaction would decrease millimoles of salt but would increase millimoles of weak base C2H5NH2.

Finally we would have

Salt millimoles=(20×0.4)-(50×0.12)=2

Weak base millimoles=(30×0.2)+(50×0.12)=12

It is basic buffer. Now for basic buffer

pOH=pKb+log(salt/weak base)

pOH=5+log(2/12)=4.1

pH=14-pOH=9.9